if f(x) = sin-1x , show that ( 1-x2) f'' (x) -x f'(x) = 0 . Hence prove that , f(n+2)(0) = n2 fn (0)


1
Expert's answer
2022-05-31T10:56:49-0400
"f(x)=\sin^{-1}(x)"

"f'(x)=\dfrac{1}{\sqrt{1-x^2}}"

"f''(x)=-\dfrac{1}{2(1-x^2)^{3/2}}(-2x)=\dfrac{x}{(1-x^2)^{3/2}}"

Then

"( 1-x^2) f''(x) -xf'(x)"

"=(1-x^2)(\dfrac{x}{(1-x^2)^{3/2}})-x(\dfrac{1}{\sqrt{1-x^2}})"

"=\dfrac{x}{\sqrt{1-x^2}}-\dfrac{x}{\sqrt{1-x^2}}=0, -1<x<1"


"( 1-x^2) f''(x) -xf'(x)=0"

"( 1-x^2) f''(x)=xf'(x)"

Differentiate with respect to "x"

"-2xf''(x)+( 1-x^2) f'''(x) =f'(x)+xf''(x)"

"( 1-x^2) f'''(x) =f'(x)+3xf''(x)"

Differentiate with respect to "x"


"( 1-x^2)f^{(4)}(x)-2xf'''(x)=f''(x)+3f''(x)"

"+3xf'''(x)"

"( 1-x^2)f^{(4)}(x)=4f''(x)+5xf'''(x)"

Differentiate with respect to "x"

"( 1-x^2)f^{(5)}(x)-2xf^{(4)}(x)=4f'''(x)+5f'''(x)"

"+5xf^{(4)}(x)"


"( 1-x^2)f^{(5)}(x)=9f'''(x)+7xf^{(4)}(x)"

Differentiate with respect to "x"

"( 1-x^2)f^{(6)}(x)-2xf^{(5)}(x)=9f^{(4)}(x)+7f^{(4)}(x)"

"+7xf^{(5)}(x)"

"( 1-x^2)f^{(6)}(x)=16f^{(4)}(x)+9xf^{(5)}(x)"

Let "P(n)" be the proposition that


"( 1-x^2)f^{(n+2)}(x)=n^2f^{(n)}(x)+(2n+1)xf^{(n+1)}(x)"

"n=1,2,..."

Basic Step

"P(1)" is true because

"( 1-x^2) f^{(1+2)}(x) =(1)^2f^{(1)}(x)+(2(1)+1)xf^{(1+1)}(x)"

Inductive Step

We assume that "P(k)" holds for an arbitrary positive integer "k"

That is, we assume that


"( 1-x^2)f^{(k+2)}(x)=k^2f^{(k)}(x)+(2k+1)xf^{(k+1)}(x)"

Differentiate with respect to "x"


"( 1-x^2)f^{(k+2+1)}(x)-2xf^{(k+2)}(x)=k^2f^{(k+1)}(x)"

"+(2k+1)f^{(k+1)}(x)+(2k+1)xf^{(k+1+1)}(x)"

Then

"( 1-x^2)f^{((k+1)+2)}(x)=(k+1)^2f^{(k+1)}(x)"

"+(2(k+1)+1)xf^{((k+1)+1)}(x)"

This last equation shows that "P(k + 1)" is true under the assumption that "P(k)" is true. This completes the inductive step.

We have completed the basis step and the inductive step, so by mathematical induction we know that "P(n)" is true for all positive integers "n." That is


"( 1-x^2)f^{(n+2)}(x)=n^2f^{(n)}(x)+(2n+1)xf^{(n+1)}(x)"

"n=1,2,..."

Substitute "x=0"


"f^{(n+2)}(0)=n^2f^{(n)}(0)"